Proof: 1+sinx/1-sinx=(tanx+secx)^2
Answer by Boreal(15235)

(tanx+secx)^2=(sin^2 x/cos^2 x)+2(sin x/cos^2 x)+(1/cos^2 x), because sec x=(1/cos x)
=over common denominator of cos^2 x, which is also 1-sin^2 x,
the numerator is sin^2 x+ 2 sin x +1=(1+sin x)^2
Therefore, the whole fraction is (1+sin x)(1+sin x)/(1+sin x)(1-sin x), because difference of squares.
The (1+sin x ) cancel,
and the answer is (1+sin x)/(1-sin x)